OpenStudy (anonymous):

b²-16b+64=16 SOLVE PL0X

OpenStudy (ash2326):

we have \[b^2-16b+64=16\] subtract both sides by 16 \[b^2-16b+48\] let's find factors of 48 ,which sums up to -48 -12 and -4 -12+(-4)=-16 -12*-4=48 so we have now \[b^2-12b-4b+48\] \[b(b-12)-4(b-12)\] we get finally \[(b-4)(b-12)\] so the roots are b=4 and b=12

OpenStudy (mimi_x3):

woops, i didnt see the 16, lol. im blind xD

OpenStudy (anonymous):

thanks